This question has two parts: What is the pH of a 0.1 M solution of HF? b) What is the percent dissociation of a 0.1 M HF solution? PLEASE SHOW WORK TYPED. It's hard to understand some handwritings so this is the easiest way I can ensure I can actually read the answer and understand it.
Answer #1

Solution:

HF is a weak acid, hence it dissociated as,

HF = H+ + F-

0.10---0-----0 (initial concentration)

(0.10- X)---X---X (at equilibrium)

Dissociation constant, Ka = X x X / (0.10 - X)

Ka for HF = 3.5 x 10^-4

Therefore, 3.5 x 10^-4 = X2 / 0.10-X

X2 + 3.5 x.10^-4 X - 3.5 x 10^-5 = 0

On solving the quadratic equation,

X = 5.74 x 10^-3 M

pH = -log X

pH = -log ( 5.74 x 10^-3)

pH = 2.24

Part B)

Percentage dissociation =( X/ 0.10 M ) x 100

= 5.74 x10^-3 x 100/ 0.10 = 5.74 %
Answer #2

Given:

  • Concentration of HF (C) = 0.1 M

  • Acid dissociation constant (Ka) of HF = 6.6 × 10⁻⁴


Part a).  pH of the Solution :-

  1. Dissociation Equation:

    HFH++F

  2. Equilibrium Expression:

    Ka=[H+][F][HF]

  3. Let x = [H⁺] at equilibrium:

    • Initial:

      [HF]=0.1M,[H+]=0,[F]=0

    • Change:

      x,+x,+x

    • Equilibrium:

      [HF]=0.1x,[H+]=x,[F]=x

  4. Substitute into Ka:

    6.6×104=xx0.1xx20.1

    (Approximation: x0.1, so 0.1x0.1)

  5. Solve for x:

    x2=6.6×105    x=6.6×1050.0081M

    (This is the [H⁺] at equilibrium.)

  6. Calculate pH:

    pH=log[H+]=log(0.0081)2.09

 Answer:
The pH of the 0.1 M HF solution is 2.09.




Part b). Percent Dissociation :-

  1. Formula:

    %Dissociation=([H+][HF]initial)×100

  2. Substitute [H⁺] = 0.0081 M and [HF]₀ = 0.1 M:

    %Dissociation=(0.00810.1)×100=8.1%

Final Answer:
The percent dissociation of the 0.1 M HF solution is 8.1%.


 


answered by: Harshwardhan kunal
Answer #3

Given Data:

  • Concentration of HF, C=0.1MC = 0.1 \, \text{M}

  • KaK_a for HF = 6.6×1046.6 \times 10^{-4}


(a) To calculate the pH:

HF is a weak acid, so it dissociates partially:

HFH++F\text{HF} \rightleftharpoons \text{H}^+ + \text{F}^-

Set up an ICE table:

SpeciesInitial (M)Change (M)Equilibrium (M)
HF0.1-x0.1 - x
H⁺0+xx
F⁻0+xx

Apply the expression for the acid dissociation constant:

Ka=[H+][F][HF]K_a = \frac{[H^+][F^-]}{[HF]}6.6×104=x20.1x6.6 \times 10^{-4} = \frac{x^2}{0.1 - x}

Since KaK_a is relatively small, we can approximate by assuming 0.1x0.10.1 - x \approx 0.1, simplifying the equation:

6.6×104=x20.16.6 \times 10^{-4} = \frac{x^2}{0.1}x2=(6.6×104)×0.1=6.6×105x^2 = (6.6 \times 10^{-4}) \times 0.1 = 6.6 \times 10^{-5}x=6.6×1058.13×103Mx = \sqrt{6.6 \times 10^{-5}} \approx 8.13 \times 10^{-3} \, \text{M}pH=log(8.13×103)2.09\text{pH} = -\log(8.13 \times 10^{-3}) \approx 2.09

pH ≈ 2.09


(b) To calculate the percent dissociation:

Percent dissociation=[H+][HF]initial×100\text{Percent dissociation} = \frac{[H^+]}{[HF]_{\text{initial}}} \times 100Percent dissociation=8.13×1030.1×1008.13%\text{Percent dissociation} = \frac{8.13 \times 10^{-3}}{0.1} \times 100 \approx 8.13\%

Percent dissociation ≈ 8.13%


answered by: Monu Kumar Gupta